OpenStudy (anonymous):

please help i am a senior who is failing and doing it on my own!!: what is the equation of the line written in slope-interept form that passes through the point (6,1) and is perpendicular to the line y+2=2(x-3)?

OpenStudy (godlovesme):

hang on i'll help u! :)

OpenStudy (anonymous):

thanks u so much i need so much help!

OpenStudy (godlovesme):

first lets write the y+2= 2(x-3) into y-intercept form distribute 2 and subtract two can u do that? :)

OpenStudy (freckles):

These are just the steps to take. But I will let god help you. My first step: Write the line in y=mx+b form My second step: Identify m. This is the slope of y=mx+b. Third step: Realize -1/m is the slope of the perpendicular line. Fourth step: Perpendicular line to y=mx+b is going to be in the form y=-1/m * x+c . Fifth step: Find c by using your point given. That is replace (x,y) with (6,1) and solve for c. -----

OpenStudy (freckles):

That sounded bad

OpenStudy (freckles):

I mean I will let godlovesme help

OpenStudy (godlovesme):

wow @freckles u r correct!! :)

OpenStudy (godlovesme):

 does that make sense @priea :)